The sum of the digits of a two digit number is 9. The number is twice the product of its digits. Find the number.
Answer by stanbon(75887)

The sum of the digits of a two digit number is 9. The number is twice the product of its digits. Find the number.
----------
Let the number be 10t+u
-----
Equations:
t + u = 9
10t+u = 2tu
------
Substitute for "t" and solve for "u":
10(9-u) + u = 2(9-u)u
------
90 - 9u = 18u - 2u^2
----
2u^2 - 27u + 90 = 0
------
(2u-15)(u-6) = 0
u = 7.5 (extraneous because the answeer must be a whole number)
u = 6
---
Solve for "t":
t + u = 9
t + 6 = 9
t = 3
----
Ans: 10t+u = 36
----------------------
Cheers,
Stan H.
------------